Brian has half of his investments in stock paying a 7% dividend and the other half in a stock paying 13% interest. If his total annual interest is $730, how much does he have invested?

Found 3 solutions by ikleyn, greenestamps, MathTherapy:
Answer by ikleyn(52908)

.
Brian has half of his investments in stock paying a 7% dividend
and the other half in a stock paying 13% interest.
If his total annual interest is $730, how much does he have invested?
~~~~~~~~~~~~~~~~

Let x be that half, which is invested at 7%.
Then the otherr half is an equal amount of money.


In a year, first half produces interest of 0.07x dollars.
The other half produces an interest of 0.13x dollars in a year.
Both halves produce the total annual 0.7x+0.13x = 0.2x.


We are given that

    0.2x = 730 dollars.


Hence,  x = 730/0.2 = 7300/2 = 3650.


Thus one half of the invested amount (each half) is 3650 dollars 
and the total invested amount is 2*3650 = 7300 dollars.    ANSWER


CHECK.  0.07*3650 + 0.13*3650 = 730 dollars, total annual interest.   ! correct !

Solved.



Answer by greenestamps(13215)



If the amounts invested at 7% and 13% are equal, then the average rate of return is 10%. So

730 = 0.10(x)
x=7300

ANSWER: $7300

CHECK: .07(3650)+.13(3650)=.20(3650)=730


Answer by MathTherapy(10557)

Brian has half of his investments in stock paying a 7% dividend and the other half in a stock paying 13% interest. If his total annual interest is $730, how much does he have invested?
Let amount invested be A
Then amount invested at 7% and 13% is .5A
We then get: .07(.5A) + .13(.5A) = 730
                   .035A + .065A = 730
                             .1A = 730
          Amount invested, or